OpenStudy (anonymous):

How do you solve 48+25=3x

OpenStudy (anonymous):

48+25=73 3x=73 x=73/3 or 24.33333

OpenStudy (anonymous):

How did you get it?

OpenStudy (anonymous):

easiest way to solve it is add 48+25 than divide both sides by 3.

OpenStudy (anonymous):

@Staystrong214 don't worry stay strong bro

OpenStudy (anonymous):

You get 24.333(repeat). So suparon is correct.

OpenStudy (anonymous):

Alright.. What about (4/5+2/3)(1/3)-x=3

OpenStudy (anonymous):

First use PEMDAS. Solve part of the equation.

OpenStudy (anonymous):

I don't know how though!

OpenStudy (anonymous):

PARENTHESES EXPONENTS MULTIPLICATION DIVISION ADDITION AND SUBTRACTION In order.

OpenStudy (anonymous):

How do you do that with decimals?